OpenStudy (anonymous):

Rename any improper answers to mixed numbers. 14 5/4 = 17 58/40 = 16 15/12 = 74 19/14 =

OpenStudy (hoblos):

14 5/4 = 15 1/4

OpenStudy (hoblos):

17 58/40 = 18 18/40 16 15/12 = 17 3/12 74 19/14 = 75 5/14